OpenStudy (anonymous):

What is the equation, in standard form, of a parabola that contains the following points?(–2, –20), (0, –4), (4, –20)

OpenStudy (anonymous):

@Hero or @Preetha can one of you help me?

OpenStudy (anonymous):

How do you do that?

OpenStudy (anonymous):

how do you solve for a? sorry, still getting the hang of this

OpenStudy (ranga):

I think you may have to start with the general equation of a parabola: y = ax^2 + bx + c Substitute the three points: (–2, –20), (0, –4), (4, –20) into the equation and solve for a, b and c.

OpenStudy (ranga):

The middle point right away gives c = -4

OpenStudy (ranga):

The vertex for this parabola happens to be at (1, -2).

OpenStudy (ranga):

I am getting y = -2x^2 + 4x - 4

OpenStudy (anonymous):

Can one of you walk me through this?

OpenStudy (ranga):

y = ax^2 + bx + c Substitute (–2, –20), (0, –4), (4, –20). First try the middle point. when x = 0, y = -4 -4 = 0 + 0 + c. c = -4 y = ax^2 + bx - 4 put x = -2, y = -20 -20 = 4a - 2b - 4 -16 = 4a - 2b -8 = 2a - b 2a - b = -8 ---- equation (1) put x = 4, y = -20 -20 = 16a + 4b - 4 -16 = 16a + 4b -4 = 4a + b 4a + b = -4 ---- equation (2) Add equations (1) and (2): 6a = -12 a = -2 Put a = -2 in equation (1): 2(-2) - b = -8 -4 - b = -8 b = 4 Therefore, y = -2x^2 + 4x - 4